# Emergent Berezinskii-Kosterlitz-Thouless deconfinement in super-Coulombic plasmas

### v1

@@ Abstract
Introduction – Berezinskii-Kosterlitz-Thouless (BKT) criticality is a paradigmatic example of a phase transition beyond Landau’s theory of symmetry breaking. The BKT universality class has been shown to underlie critical phenomena in numerous physical systems displaying superfluidity and superconductivity in thin films [1, 2], Josephson junction arrays [3, 4], ultracold gases [5, 6], polaritonic systems [7, 8], among others.
@@
The underlying physical mechanism of the BKT transition is a proliferation of vortices, acting as disorder operators in the statistical mechanics of two-dimensional $U(1)$ symmetric models [9, 10]. The transition is then driven by the subtle energy-entropy competition between the energy cost of unbinding opposite charge vortex pairs and the configurational entropy gain. Crucially, both follow logarithmic scaling with respect to the spatial separation between vortex anti-vortex pairs. The resulting phase diagram, when described in terms of a vortex Coulomb gas, features a low-temperature phase with confined, tightly bound vortex pairs and a high-temperature phase characterized by deconfined, free vortices.
@@
Figure 1: Monte Carlo snapshots of a super-Coulombic gas with confining power-law, $r^{\sigma}$, interaction, with $\sigma=0.25$ and particle density $\rho=0.01$ in its (a) confined and (b) deconfined phases separated by a BKT phase transition. The orange (blue) circles indicate positive (negative charges). (c) The power law vs temperature phase diagram for fixed particle density $\rho=0.01$. (d) The particle density vs temperature phase diagram for fixed power law $\sigma=0.25$. The red points indicate BKT critical points observed using our simulations; the phase boundary is constructed by interpolating between them.
@@
Recent experimental interest in systems with arbitrary long-range interactions [11, 12, 13, 14, 15] has posed questions about the fate of the BKT transition in the presence of a more general power-law interaction, $r^{\sigma}$ [16, 17, 18, 19, 20, 21]. In particular, two recent works [19, 18], studying the long-wavelength physics of super-Coulombic gases, i.e., gases with longer-ranged than Coulomb interactions, proposed, by means of Renormalization Group (RG) and dielectric screening (with a length-scale dependent dielectric function) calculations, that generic super-Coulombic gases universally display a screening crossover to an effective Coulomb description. This emergence is also expected in higher dimensions [22, 18].
@@
An intriguing corollary of the aforementioned Coulomb emergence is that two-dimensional super-Coulombic gases were predicted [18, 19] to display a BKT class confinement-deconfinement transition at finite temperatures. This prediction starkly contrasts with the expectation of confinement at all temperatures, based on a naive energy-entropy argument that neglects screening.
@@
In this Letter, we employ numerically exact calculations to showcase the super-Coulombic to Coulomb crossover and the subsequent BKT confinement-deconfinement transition in a concrete microscopic model. We identify low (high) temperature confining (deconfining) phases, see Figs. 1a and 1b. To probe the emergent Coulomb interaction at long distances, we introduce a numerical method for extracting the interaction between test charges and determining the associated RG flow. The effective Coulomb coupling displays the universal Nelson-Kosterlitz jump [23] at the deconfinement transition point, as expected for the BKT universality. The emergent Coulomb coupling in the bound phase increases sharply at low temperatures, reflecting power-law screening [18], unlike conventional Coulomb gases, where it saturates to its bare value. Lastly, by studying this problem for different power laws and densities, we establish the generality of Coulomb emergence across a wide range of microscopic parameters and interaction forms, as shown in Figs. 1c and 1d.
@@
The aforementioned argument is incorrect, as it neglects the predicted screening effects [18, 19] resulting in a super-Coulombic to Coulomb crossover, as will be demonstrated explicitly in this work.
Effective Coulomb coupling – The standard technique for studying the BKT transition in Coulomb gases involves tracking the effective Coulomb coupling as an order parameter. Typically, the renormalization of the bare Coulomb coupling is extracted in terms of the charge density-density correlation function [24, 25]. However, for a super-Coulombic gas displaying Coulomb emergence, the bare Coulomb coupling or the charge densities of emergent Coulomb particles are hard to define precisely, rendering the standard approach inapplicable. To address this, we devise a method to directly extract the effective potential experienced by fixed test charges in the presence of a super-Coulombic plasma medium.
@@
For the case of Coulomb emergence [18, 19], we anticipate an asymptotic Coulomb-like form $e^{-\beta V_{\text{eff}}(r)}=e^{-\beta(\kappa n^{2}_{t}\log(r)+c)}$, where $c$ is a constant, $n_{t}$ is the magnitude of the test charges and $\kappa$ is the long wavelength effective Coulomb constant. For intermediate scales, we define a running Coulomb coupling $\kappa(r)$ by measuring $V_{\text{eff}}(r)$ at two reference points, $r$ and $r/\sqrt{2}$:
$\kappa(r)\equiv-\frac{1}{\beta n_{t}^{2}}\log\left(\frac{e^{-\beta V_{\text{eff}}(r)}}{e^{-\beta V_{\text{eff}}\left(\frac{r}{\sqrt{2}}\right)}}\right)/\log(\sqrt{2}).$ (3)
For a super-Coulombic gas displaying emergent Coulomb behavior, $\lim_{r\to\infty}\kappa(r)$ should either saturate to a constant $\kappa_{\infty}$ in the bound phase with $\kappa_{\infty}>4T$ or vanish in the unbound phase of a Coulomb plasma. The transition between the phases should occur at a temperature satisfying the universal BKT relation $T=\kappa_{\infty}/4$ [23]. This behaviour contrasts with bare super-Coulombic potentials, where $\lim_{r\to\infty}\kappa(r)\sim r^{\sigma}/\log(r)$ diverges to infinity.
For our results, we perform Markov Chain Monte Carlo simulations [25, 26] of a canonical ensemble of neutral 2D super-Coulombic plasmas with $N$ charges at fixed particle density $\rho$, implicitly fixing a simulation box size $L$. To avoid boundary effects, the potential is measured in the bulk of the box, with a “boundary buffer” of length $L/3$. For our finite-sized system, we define the long-wavelength Coulomb coupling as $\kappa_{L}\equiv\kappa(L/3)$. All energies and distances are measured in units of the bare coupling $K$ and cut-off $a$, respectively. Additional details about the simulation and an explanation of how our observable, defined in Eq. 2, is computed are presented in Ref. 24.
@@
Results – In what follows, we present results for a $\sigma=0.25$, $\rho=0.01$ super-Coulombic gas unless specified otherwise. We begin by examining the functional form of $V_{\text{eff}}(r)$. By way of example, in Fig. 2a, we plot $V_{\text{eff}}(r)$, contrasting it against the bare potential and an asymptotic Coulomb-like potential. We see that as $r$ increases, $V_{\text{eff}}(r)$ very quickly crosses over from its bare power-law form to a logarithmic Coulomb-like curve. We note that this is strong evidence for Coulomb emergence [18]. Moreover, in Fig. 2b, plotting $V_{\text{eff}}(r)$ for different temperatures, we see that the effective coupling decreases sharply with increasing temperature, reminiscent of Coulomb gas like behavior in the vicinity of a BKT confinement-deconfinement transition.
@@
The transition between the two aforementioned phases is marked by a crossing point between the different finite system $\kappa_{L}$ vs $T$ curves, implying the existence of scale-invariant physics at this point. Remarkably, this point also lies on the $\kappa=4T$ universal BKT line, certifying its universality class. The thermodynamic extrapolation of crossing points between finite system curves determines $T_{c}$, as quoted above. In Ref. 24, we showcase this transition for different $\sigma$ and $\rho$ parameters, attesting that the $\kappa=4T$ relation holds across power-laws and densities, demonstrating the predicted universal occurrence of BKT transitions in generic super-Coulombic gases.
@@
Fig. 3a shows that in the confined phase $\kappa_{L}$ gradually increases with the linear box size $L$, in contrast to the expected constant Coulomb coupling $\kappa$ in the thermodynamic limit. We attribute this to a finite-size RG flow associated with the super-Coulombic to Coulomb crossover, that saturates only at length scales greater than $L$. To support our claim, we study the scaling of $\kappa_{L}$ vs $1/L$ for different temperatures in the confined phase in Fig. 4a. To quantify the thermodynamic convergence of $\kappa_{L}$ more accurately, we extract $\kappa_{\infty}$ by fitting our results to the ansatz $\kappa_{L}=\kappa_{\infty}-AL^{-\omega}$, where $A$ and $\omega$ are fitting parameters. The extrapolation as depicted in Fig. 4a, shows that $\kappa_{L}$ indeed approaches a constant value $\kappa_{\infty}$ in the limit $L\to\infty$, suggesting Coulomb emergence. This convergence slows down at lower temperatures, as the screening length grows [18], requiring even larger system sizes for $\kappa_{L}$ to saturate.
The temperature scaling of $\kappa_{\infty}$ in Fig. 4a suggests a monotonic divergence of the effective Coulomb coupling with a decreasing temperature, deeper in the confining phase. This should be contrasted with the standard pure Coulomb case, which is expected to saturate to its bare value at low temperatures, as schematically depicted in Fig. 4b. This behavior is consistent with the RG calculations in Ref. 18, which predict an exponential divergence of $\kappa_{\infty}$ at lower temperatures.
Since $\kappa_{L}$ flows for finite-sized systems, the length-scale at which the Coulomb regime fully settles is captured only when $\kappa_{L}\sim\kappa_{\infty}$. Nevertheless, it is fruitful to study the temperature dependence of the length-scale for which the crossover emerges, i.e., when $V_{\text{eff}}$ displays significant deviation from its bare form. In Fig. 2, we observe that for the specific choice of parameters, $V_{\text{eff}}$ departs from the microscopic super-Coulombic potential already at scales comparable to the microscopic cutoff. To resolve the evolution clearly, in Ref. 24, we study $V_{\text{eff}}$ at extremely low density, for which there is a separation of scales of the form $a\ll\xi_{c}\ll\rho^{-1/2}$. From this analysis, we observe that this length increases with decreasing temperature.
Alternatively, Ref. 18 defines $\xi_{c}$ as the length-scale where the renormalized super-Coulombic contribution in $V_{\text{eff}}$ becomes comparable in magnitude to Coulomb contribution generated under coarse-graining RG. Up to an $O(1)$ constant, this can be extracted using the relation $\xi_{c}\times O(1)=\kappa_{\infty}^{1/\sigma}$ [18]. In Ref. 24, we show that this length-scale diverges at low temperatures, similarly to $\kappa_{\infty}.$
@@
It is interesting to extend our study to driving the BKT transition by varying the power law $\sigma$ at a fixed temperature and density. In Fig. 5, fixing $T=0.078$ and $\rho=0.01$, we observe a deconfined phase with $\kappa=0$ for $\sigma\lesssim 0.25$, undergoing a transition to a confined phase with $\kappa\neq 0$ at a critical point $\sigma=0.25$, respecting the relation $\kappa=4T$. The phase diagram corroborates the intuitive expectation that gases with larger power laws bind charges more strongly, thus requiring higher temperatures to deconfine. Consequently, $T_{\text{BKT}}$ should increase with increasing $\sigma$. The BKT crossings shown in [24] for different power laws exemplify this behavior.
Similarly, we investigate the phase diagram as a function of particle density $\rho$ for a fixed power law $\sigma$ and temperature $T$. It is known that in Coulomb gases, increasing the particle density decreases $T_{\text{BKT}}$ [25]. In Fig. 5b, we show that super-Coulombic gases display similar behavior. Fixing $\sigma$ to 0.25 and $T$ to 0.078, we observe that as $\rho$ is increased, the gas undergoes a phase transition from a confined to a deconfined phase. Moreover, in [24] we show $\kappa_{L}$ vs $T$ BKT crossings with $\sigma=0.25$ for varying densities display an inverse relationship between $\rho$ and $T_{\text{BKT}}$. Combining these results, we construct two-dimensional confinement-deconfinement phase diagrams in the $\sigma$ vs $T$ (Fig. 1c) and $\rho$ vs $T$ (Fig. 1d) planes.
@@
Summary and discussion – Our results present direct and multifaceted numerical confirmation of novel dielectric screening in a two-dimensional super-Coulombic gas, showing the emergence of an effective Coulomb interaction beyond the screening length, corroborating recent RG predictions [18, 19]. This emergent behavior arises despite the underlying microscopic power-law interaction being longer-ranged than Coulomb, and is reflected in the asymptotic form of the effective potential beyond the crossover length scale $\xi_{c}$, and the saturation of the renormalized Coulomb coupling $\kappa_{\infty}$ at large distances.
@@
A notable observation is that the crossover into a true Coulomb description is long-tailed, saturating at distances much longer than length-scales where screening becomes significant. We defer a detailed study of this intermediate crossover regime to future work. Additionally, the predicted anomalous screening behavior in the deconfined phase, characterized by power-law rather than exponential (Debye-Huckel) decay of interactions in Ref. 18 remains an intriguing open direction for theoretical and numerical study. Finally, generalizing these ideas to higher dimensions, where Coulomb emergence is still expected [22, 18], presents a natural extension of this work.

### v2

@@ Abstract
Introduction – Berezinskii-Kosterlitz-Thouless (BKT) criticality is a paradigmatic example of a phase transition beyond Landau’s theory of symmetry breaking. The BKT universality class has been shown to underlie critical phenomena in numerous physical systems displaying superfluidity and superconductivity in thin films [Bishop_1979, Bishop_1980], Josephson junction arrays [Resnick_1981, Newrock_2000], ultracold gases [Hadzibabic_2006, Murthy_2015], polaritonic systems [Nitsche_2014, Caputo_2018], among others.
@@
The underlying physical mechanism of the BKT transition is a proliferation of vortices, acting as disorder operators in the statistical mechanics of two-dimensional $U(1)$ symmetric models [Kosterlitz_1973, Jose_1977]. The transition is then driven by the subtle energy-entropy competition between the energy cost of unbinding opposite charge vortex pairs and the configurational entropy gain. Crucially, both follow logarithmic scaling with respect to the spatial separation between vortex anti-vortex pairs. The resulting phase diagram, when described in terms of a vortex Coulomb gas, features a low-temperature phase with confined, tightly bound vortex pairs and a high-temperature phase characterized by deconfined, free vortices.
@@
Figure 1: Monte Carlo snapshots of a super-Coulombic gas with confining power-law, $r^{\sigma}$, interaction, with $\sigma=0.25$ and particle density $\rho=0.01$ in its (a) confined and (b) deconfined phases separated by a BKT phase transition. The orange (blue) circles indicate positive (negative) charges. (c) The power law vs temperature phase diagram for fixed particle density $\rho=0.01$. (d) The particle density vs temperature phase diagram for fixed power law $\sigma=0.25$. The red points indicate BKT critical points observed using our simulations; the phase boundary is constructed by interpolating between them.
@@
Recent experimental interest in systems with arbitrary long-range interactions [Yan_2013, Bruzewicz_2019, Bendkowsky_2009, Dunning_2024, DeSalvo_2015] has posed questions about the fate of the BKT transition in the presence of a more general power-law interaction, $r^{\sigma}$ [Giachetti_2021, Defenu_2023, Radzihovsky_2024, Giachetti_2023, Yao_2024, Xiao_2024]. In particular, two recent works [Giachetti_2023, Radzihovsky_2024], studying the long-wavelength physics of super-Coulombic gases, i.e., gases with longer-ranged than Coulomb interactions, proposed, by means of Renormalization Group (RG) and dielectric screening (with a length-scale dependent dielectric function) calculations, that generic super-Coulombic gases universally display a screening crossover to an effective Coulomb description. This emergence is also expected in higher dimensions [Herbut_2003, Radzihovsky_2024].
@@
An intriguing corollary of the aforementioned Coulomb emergence is that two-dimensional super-Coulombic gases were predicted [Radzihovsky_2024, Giachetti_2023] to display a BKT class confinement-deconfinement transition at finite temperatures. This prediction starkly contrasts with the expectation of confinement at all temperatures, based on a naive energy-entropy argument that neglects screening.
@@
In this Letter, we employ numerically exact calculations to showcase the super-Coulombic to Coulomb crossover and the subsequent BKT confinement-deconfinement transition in a concrete microscopic model. We identify low (high) temperature confining (deconfining) phases, see Figs. 1a and 1b. To probe the emergent Coulomb interaction at long distances, we introduce a numerical method for extracting the interaction between test charges and determining the associated RG flow. The effective Coulomb coupling displays the universal Nelson-Kosterlitz jump [Nelson_1977] at the deconfinement transition point, as expected for the BKT universality. The emergent Coulomb coupling in the bound phase increases sharply at low temperatures, reflecting power-law screening [Radzihovsky_2024], unlike conventional Coulomb gases, where it saturates to its bare value. Lastly, by studying this problem for different power laws and densities, we establish the generality of Coulomb emergence across a wide range of microscopic parameters and interaction forms, as shown in Figs. 1c and 1d.
@@
The aforementioned argument is incorrect, as it neglects the predicted screening effects [Radzihovsky_2024, Giachetti_2023] resulting in a super-Coulombic to Coulomb crossover, as will be demonstrated explicitly in this work.
Effective Coulomb coupling – The standard technique for studying the BKT transition in Coulomb gases involves tracking the effective Coulomb coupling as an order parameter. Typically, the renormalization of the bare Coulomb coupling is extracted in terms of the charge density-density correlation function [SM, Orkoulas_1996]. However, for a super-Coulombic gas displaying Coulomb emergence, the bare Coulomb coupling or the charge densities of emergent Coulomb particles are hard to define precisely, rendering the standard approach inapplicable. To address this, we devise a method to directly extract the effective potential experienced by fixed test charges in the presence of a super-Coulombic plasma medium.
@@
For the case of Coulomb emergence [Radzihovsky_2024, Giachetti_2023], we anticipate an asymptotic Coulomb-like form $e^{-\beta V_{\text{eff}}(r)}=e^{-\beta(\kappa n^{2}_{t}\log(r)+c)}$, where $c$ is a constant, $n_{t}$ is the magnitude of the test charges and $\kappa$ is the long wavelength effective Coulomb constant. For intermediate scales, we define a running Coulomb coupling $\kappa(r)$ by measuring $V_{\text{eff}}(r)$ at two reference points, $r$ and $r/\sqrt{2}$:
$\kappa(r)\equiv\frac{1}{n^{2}_{t}}\left(\frac{V_{\text{eff}}(r)-V_{\text{eff}}\left(\frac{r}{\sqrt{2}}\right)}{\log(\sqrt{2})}\right)$ (3)
For a super-Coulombic gas displaying emergent Coulomb behavior, $\lim_{r\to\infty}\kappa(r)$ should either saturate to a constant $\kappa_{\infty}$ in the bound phase with $\kappa_{\infty}>4T$ or vanish in the unbound phase of a Coulomb plasma. The transition between the phases should occur at a temperature satisfying the universal BKT relation $T=\kappa_{\infty}/4$ [Nelson_1977]. This behaviour contrasts with bare super-Coulombic potentials, where $\lim_{r\to\infty}\kappa(r)\sim r^{\sigma}$ diverges to infinity.
For our results, we perform Markov Chain Monte Carlo simulations [Orkoulas_1996, Saito_1981] of a canonical ensemble of neutral 2D super-Coulombic plasmas with $N$ charges at fixed particle density $\rho$, implicitly fixing a simulation box size $L$. To avoid boundary effects, the potential is measured in the bulk of the box, with a “boundary buffer” of length $L/3$. For our finite-sized system, we define the long-wavelength Coulomb coupling as $\kappa_{L}\equiv\kappa(L/3)$. All energies and distances are measured in units of the bare coupling $K$ and cut-off $a$, respectively. Additional details about the simulation and an explanation of how our observable, defined in Eq. 2, is computed are presented in the Supplemental Materials.
@@
Results – In what follows, we present results for a $\sigma=0.25$, $\rho=0.01$ super-Coulombic gas unless specified otherwise. We begin by examining the functional form of $V_{\text{eff}}(r)$. By way of example, in Fig. 2a, we plot $V_{\text{eff}}(r)$ at a fixed temperature, contrasting it against the bare potential and an asymptotic Coulomb-like potential. We see that as $r$ increases, $V_{\text{eff}}(r)$ very quickly crosses over from its bare power-law form to a logarithmic Coulomb-like curve. We note that this is strong evidence for Coulomb emergence [Radzihovsky_2024]. Moreover, in Fig. 2b, plotting $V_{\text{eff}}(r)$ for different temperatures, we see that the effective coupling decreases sharply with increasing temperature, reminiscent of Coulomb gas like behavior in the vicinity of a BKT confinement-deconfinement transition.
@@
The transition between the two aforementioned phases is marked by a crossing point between the different finite system $\kappa_{L}$ vs $T$ curves, implying the existence of scale-invariant physics at this point. Remarkably, this point also lies on the $\kappa=4T$ universal BKT line, certifying its universality class. The thermodynamic extrapolation of crossing points between finite system curves determines $T_{c}$, as quoted above. In the supplementary [SM], we showcase this transition for different $\sigma$ and $\rho$ parameters, attesting that the $\kappa=4T$ relation holds across power-laws and densities, demonstrating the predicted universal occurrence of BKT transitions in generic super-Coulombic gases.
@@
Fig. 3a shows that in the confined phase $\kappa_{L}$ gradually increases with the linear box size $L$, in contrast to the expected constant Coulomb coupling $\kappa$ in the thermodynamic limit. We attribute this to a finite-size RG flow associated with the super-Coulombic to Coulomb crossover, that saturates only at length scales greater than $L$. To support our claim, we study the scaling of $\kappa_{L}$ vs $1/L$ for different temperatures in the confined phase in Fig. 4a. To quantify the thermodynamic convergence of $\kappa_{L}$ more accurately, we extract $\kappa_{\infty}$ by fitting our results to the ansatz $\kappa_{L}=\kappa_{\infty}-AL^{-\omega}$, where $A$ and $\omega$ are fitting parameters. The extrapolation as depicted in Fig. 4a, shows that $\kappa_{L}$ indeed approaches a constant value $\kappa_{\infty}$ in the limit $L\to\infty$, suggesting Coulomb emergence. This convergence slows down at lower temperatures, as the screening length grows [Radzihovsky_2024], requiring even larger system sizes for $\kappa_{L}$ to saturate.
The temperature scaling of $\kappa_{\infty}$ in Fig. 4a suggests a monotonic divergence of the effective Coulomb coupling with a decreasing temperature, deeper in the confining phase. This should be contrasted with the standard pure Coulomb case, which is expected to saturate to its bare value at low temperatures, as schematically depicted in Fig. 4b. This behavior is consistent with the RG calculations in Ref. Radzihovsky_2024, which predict an exponential divergence of $\kappa_{\infty}$ at lower temperatures.
Since $\kappa_{L}$ flows for finite-sized systems, the length-scale at which the Coulomb regime fully settles is captured only when $\kappa_{L}\sim\kappa_{\infty}$. Nevertheless, it is fruitful to study the temperature dependence of the length-scale for which the crossover emerges, i.e., when $V_{\text{eff}}$ displays significant deviation from its bare form. In Fig. 2, we observe that for the specific choice of parameters, $V_{\text{eff}}$ departs from the microscopic super-Coulombic potential already at scales comparable to the microscopic cutoff. To resolve the evolution clearly, in Ref. SM, we study $V_{\text{eff}}$ at extremely low density, for which there is a separation of scales of the form $a\ll\xi_{c}\ll\rho^{-1/2}$. From this analysis, we observe that this length increases with decreasing temperature.
Alternatively, Ref. Radzihovsky_2024 defines $\xi_{c}$ as the length-scale where the renormalized super-Coulombic contribution in $V_{\text{eff}}$ becomes comparable in magnitude to the Coulomb contribution generated under coarse-graining RG. Up to an $O(1)$ constant, this can be extracted using the relation $\xi_{c}\times O(1)=\kappa_{\infty}^{1/\sigma}$ [Radzihovsky_2024]. In Ref. SM, we show that this length-scale diverges at low temperatures, similarly to $\kappa_{\infty}.$
@@
It is interesting to extend our study to driving the BKT transition by varying the power law $\sigma$ at a fixed temperature and density. In Fig. 5, fixing $T=0.078$ and $\rho=0.01$, we observe a deconfined phase with $\kappa=0$ for $\sigma\lesssim 0.25$, undergoing a transition to a confined phase with $\kappa\neq 0$ at a critical point $\sigma=0.25$, respecting the relation $\kappa=4T$. The phase diagram corroborates the intuitive expectation that gases with larger power laws bind charges more strongly, thus requiring higher temperatures to deconfine. Consequently, $T_{\text{BKT}}$ should increase with increasing $\sigma$. The BKT crossings shown in Ref. SM for different power laws exemplify this behavior.
Similarly, we investigate the phase diagram as a function of particle density $\rho$ for a fixed power law $\sigma$ and temperature $T$. It is known that in Coulomb gases, increasing the particle density decreases $T_{\text{BKT}}$ [Orkoulas_1996]. In Fig. 5b, we show that super-Coulombic gases display similar behavior. Fixing $\sigma$ to 0.25 and $T$ to 0.078, we observe that as $\rho$ is increased, the gas undergoes a phase transition from a confined to a deconfined phase. Moreover, in Ref. SM we show $\kappa_{L}$ vs $T$ BKT crossings with $\sigma=0.25$ for varying densities display an inverse relationship between $\rho$ and $T_{\text{BKT}}$. Combining these results, we construct two-dimensional confinement-deconfinement phase diagrams in the $\sigma$ vs $T$ (Fig. 1c) and $\rho$ vs $T$ (Fig. 1d) planes.
@@
Summary and discussion – Our results present direct and multifaceted numerical confirmation of novel dielectric screening in a two-dimensional super-Coulombic gas, showing the emergence of an effective Coulomb interaction beyond the screening length, corroborating recent RG predictions [Radzihovsky_2024, Giachetti_2023]. This emergent behavior arises despite the underlying microscopic power-law interaction being longer-ranged than Coulomb, and is reflected in the asymptotic form of the effective potential beyond the crossover length scale $\xi_{c}$, and the saturation of the renormalized Coulomb coupling $\kappa_{\infty}$ at large distances.
@@
A notable observation is that the crossover into a true Coulomb description is long-tailed, saturating at distances much longer than length-scales where screening becomes significant. We defer a detailed study of this intermediate crossover regime to future work. Additionally, the predicted anomalous screening behavior in the deconfined phase, characterized by power-law rather than exponential (Debye-Hückel) decay of interactions in Ref. Radzihovsky_2024 remains an intriguing open direction for theoretical and numerical study. Finally, generalizing these ideas to higher dimensions, where Coulomb emergence is still expected [Herbut_2003, Radzihovsky_2024], presents a natural extension of this work.